25cm^3 of magnesium Chloride compleatly reacted with 12.5cm^3 of 0.2M silver nitrate to precipitate silver Chloride .Haw many grams of magnesium Chloride are contained in 500cm^3 of the solution

1
Expert's answer
2022-03-18T16:39:03-0400

 

Moles of silver nitrate:

V = 12.5 cm3 = 12.5 mL = 0.0125 L

0.0125 L x 0.2 M = 0.0025 mol

MgCl2 + 2AgNO3 = Mg(NO3)2 + 2AgCl

The ratio of the reactants is 1 : 2

So, moles of MgCl2: 0.0025 x 1/2 = 0.00125 mol – in 25 cm3 of the solution

In 500 cm3: 0.00125 x 500 / 25 = 0.025 mol

Molar mass of MgCl2: 95.2 g/mol

Mass = 0.025 x 95.2 = 2.38 g